Reports 1-1 of 1 Clear search Modify search
Environmental Monitoring (Environmental Monitoring)
fiori, karathanasis, romero, tringali - 22:24 Monday 07 October 2019 (47122) Print this report
Task ENV02: NEB acoustic injections and tappings

On Sunday Oct 6 we tested the NEB in a way similar to what done at WEB, namely with acoustic injections and tappings on various parts and the vacuum chamber. The attached log file contains times information and details.


Using the setup shown in Figure 1 we injected white noise first in the 100Hz to 2kHz range, then in the 40-100Hz range with various levels (see log file for details). See Figure 2. The effect we saw looks smaller than at WEB (46708), but we did produce some excitation at 39.x, 45.6Hz and around 48Hz. We still have to check if something at these frequency exists in Hrec also in quiet condition. See Figure 3.


We tapped similar locations as we did for WEB (47107): Costolone, SWEB chamber, SWEB link pipe, Tower flange (E), CryoTrap bellow. We saw no effect in Hrec. Figure 4 shows the tapping on NE costolone, to be compared with the similar action at WE (link WE figure).


We attached one accelerometer to the tower vacuum chamber at second floor level (similar to WE) named ENV_WE_TOP_ACC_X. We knocked on the chamber from the second floor (as we did at WE) and doing so we excited vibration peaks at 6.25Hz, 41.4Hz, 43.8Hz, 58.4Hz, 71.2Hz (see Figure number 5) which are also visible in the accelerometer ENV_WE_ACC_Z placed on the N flange. The interferometer was not locked at the time of this measurement. We have to check if these frequencies match some existing peak in Hrec, for example the 71.2Hz.

Images attached to this report
Non-image files attached to this report
Search Help